# A Unified Variational Principle for Branching Transport Networks: Wave Impedance, Viscous Flow, and Tissue Metabolism

### 16 Mar 2026 v1

@@ 2.1 Branching law and impedance scaling
We consider a symmetric bifurcation ($N=2$) in which a parent branch of diameter $d_{p}$ divides into $N$ daughters of diameter $d_{c}$:
@@
The binary topology ($N=2$) is selected by a two-tier argument. At the static level, the topological bounding theorem of the companion paper [10] proves that the sub-linear wall cost ($p<1$) bounds the optimal branching number to small finite integers by penalizing junction tissue volume. At the dynamic level, Theorem 1 of this paper proves that among all regular fractal trees perfusing a fixed terminal population, $N=2$ uniquely maximizes the network stiffness ratio $\kappa_{\mathrm{eff}}(N)$, and consequently uniquely maximizes the minimax branching exponent $\alpha^{*}$. The two arguments are complementary: the static mechanism provides the necessary condition (small $N$), and the dynamic mechanism provides the sufficient condition (unique optimum). This establishes binary bifurcation not as an assumption but as a derived property of the unified wave-transport framework. The impedance of a cylindrical conduit scales as $Z\propto d^{-\alpha_{w}}$, where $\alpha_{w}$ depends on the wave physics. Table 1 summarizes the principal cases.
@@ 2.5 Topological selection of binary branching
where $r_{g}^{*}=r^{*}(Q_{0}/N^{g})$ are the generation-by-generation optimal radii. This decomposition holds to $\mathcal{O}(10^{-2})$: since $\alpha_{\mathrm{local}}$ varies only across $[2.898,2.907]$, the first-order gradient term $\Phi^{\prime}(r_{g}^{*})(\alpha_{t}-\alpha_{\mathrm{local}})$ is a negligible correction, leaving the Hessian $\Phi^{\prime\prime}(r_{g}^{*})$ as the dominant contribution. Consequently, $\kappa_{\mathrm{eff}}(N)=k_{t}^{\mathrm{net}}(N)/k_{w}^{\mathrm{net}}(N)$ is a strictly decreasing function of $N$, uniquely maximized by binary bifurcation ($N=2$).
@@
Numerical evaluation for porcine coronary parameters ($M=2^{11}$ terminals, full three-term cost of Eq. (4)) gives the topological hierarchy shown in Table 4. The topological selection operates not by exclusion—all values $N\leq 6$ fall within the morphometric range $\alpha_{\mathrm{exp}}=2.70\pm 0.20$—but by robustness maximization. Binary bifurcation ($N=2$) uniquely places the minimax optimum at the upper end of the distribution, maximizing the separation from the wave attractor $\alpha_{w}=2.115$ and thereby achieving the greatest possible margin of transport efficiency. The stiffness ratio $\kappa_{\mathrm{eff}}$ decreases strictly with $N$ (from $4.98$ at $N=2$ to $1.35$ at $N=6$), monotonically reducing the wave-transport equilibrium separation. The universal biological preference for binary branching is therefore a rigorous consequence of dynamic robustness optimization: $N=2$ is the unique topology that maximizes $\kappa_{\mathrm{eff}}$, and hence the unique topology that keeps the wave-transport minimax equilibrium as far as possible from the wave-dominated limit. However, the shallow sensitivity of $\alpha^{*}$ to the branching number (varying by only $\sim\!1\%$ between $N=2$ and $N=4$ in Table 4) implies a quasi-degenerate thermodynamic landscape. Rather than a model weakness, this shallow gradient explains the biological reality that occasional trifurcations ($N=3$) are not catastrophic failures of the variational principle, but sub-optimal yet accessible morphological solutions lying only slightly above the global $\kappa_{\text{eff}}$-robustness maximum.
@@ 3.2 Evolutionary Selection via Worst-Case Robustness
The optimal duty cycle $\eta^{*}$ determined by Eq. (18) represents an architectural constraint of the network architecture. Since the transport gradient is $5$ times steeper than the wave gradient at the saddle point ($|g^{\prime}|/f^{\prime}\approx 5$), the value $\eta^{*}=5/(5+1)\approx 0.833$ is strictly required for the zero-sum game to reach equilibrium. This identifies the duty cycle not as a physiological variable, but as a structural invariant of binary cardiovascular networks.
@@
The wave cost $f(\alpha)=1-(1-|\Gamma(\alpha)|^{2})^{G}$ and its derivative $f^{\prime}(\alpha^{*})$ depend only on $\alpha_{w}$, $G$, and $N$ (Eq. (2)). The transport cost $g(\alpha)$ defined by Eq. (6) is a ratio of sums of $\Phi$; its derivative $g^{\prime}(\alpha^{*})$ is likewise a ratio in which the metabolic prefactors ($b$, $m_{w}$, $\mu$) appear identically in every numerator and denominator term and therefore cancel exactly. The residual dependence on $Q_{0}$ and $\ell_{0}$ also cancels by the intensivity established in Proposition 7. Hence $\eta^{*}=R/(1+R)$ with $R=|g^{\prime}|/f^{\prime}$ depends only on $(G,p,\alpha_{w})$, confirming the claim. The sensitivity table (Table 7) provides numerical confirmation: all metabolic parameters yield $|S_{x}|<0.01$ for $\alpha^{*}$ and therefore also for $\eta^{*}$. ∎
@@ 4.1 Numerical inputs
As shown in Table 7, the minimax prediction $\alpha^{*}=2.72$ exhibits remarkable structural robustness. The exponent is essentially invariant ($|S_{x}|<0.01$) to metabolic ’noise’ such as absolute blood viscosity, proximal flow rate, or specific wall metabolic rates. It is determined strictly by the network topology ($G$) and the histological scaling ($p$), confirming it as a fundamental architectural property rather than a physiological tuning. Notably, the prediction is structurally independent of the absolute magnitude of the wall pre-factor $c_{0}$, as the scaling penalty depends only on the relative power $p$.
@@ 4.2 Central result
The Lagrange multiplier of the equal-cost constraint evaluates to $\eta^{*}=0.833$, implying that at the saddle point the transport cost gradient is steeper than the wave cost gradient (ratio $5:1$). This apparent inversion—where the static metabolic gradient outweighs the dynamic wave sensitivity—arises from the super-linearity of the radius deviation. When the global branching exponent $\alpha$ deviates from the morphometric optimum, the metabolic cost of the millions of distal vessels explodes faster than the junction-level wave reflection. This ratio implies that the coronary vasculature sits at an architecture where the transport cost is locally $5$ times more sensitive to $\alpha$ perturbations than wave costs. The distinguishing structural property of binary branching is precisely that $N=2$ maximizes the network stiffness ratio $\kappa_{\mathrm{eff}}$, providing the greatest separation between these competing gradients. As a plausibility check, the pulsatility index of coronary flow ($\mathrm{PI}=(Q_{\max}-Q_{\min})/Q_{\mathrm{mean}}\approx 2$44 [13]) independently confirms a strongly pulsatile regime in which wave-reflection costs dominate over cycle-averaged transport costs, consistent with $\eta^{*}>0.5$.
@@ 6.3 Testable predictions
The $40\%$ variation of $k_{t}^{\mathrm{net}}$ across $N\in[2,6]$ (Table 4) arises from a competition between two opposing effects: (a) the tree becomes shallower as $N$ grows, reducing the total number of summation terms and suppressing the depth-squared weighting; (b) the distal vessels in a high-$N$ tree carry lower flow, placing them in the regime where $\alpha^{*}(Q)\to(5+p)/2$ (Corollary 7 of the companion paper [10]), where the transport curvature $\Phi^{\prime\prime}(r^{*})$ is sharper. These two effects partially cancel, producing a quasi-constant $k_{t}^{\mathrm{net}}$. The net result is that the topological competition is dominated by the wave-cost curvature $k_{w}^{\mathrm{net}}\propto\ln N$, giving $\kappa_{\mathrm{eff}}(N)\propto(\ln N)^{-2.3+1}$, a strictly decreasing function of $N$.

### 17 Mar 2026 v2

@@ 2.1 Branching law and impedance scaling
We consider a symmetric bifurcation ($N=2$) in which a parent branch of diameter $d_{p}$ divides into NN daughters of diameter $d_{c}$:
@@
The binary topology ($N=2$) is selected by a two-tier argument. At the static level, the topological bounding theorem of the companion paper [10] proves that the sub-linear wall cost ($p<1$) bounds the optimal branching number to small finite integers by penalizing junction tissue volume. At the dynamic level, Theorem 1 of this paper proves that among all regular fractal trees perfusing a fixed terminal population, $N=2$ uniquely maximizes the network stiffness ratio $\kappa_{\mathrm{eff}}(N)$, and consequently uniquely maximizes the minimax branching exponent $\alpha^{*}$. The two arguments are complementary: the static mechanism provides the necessary condition (small NN), and the dynamic mechanism provides the sufficient condition (unique optimum). This establishes binary bifurcation not as an assumption but as a derived property of the unified wave-transport framework. The impedance of a cylindrical conduit scales as $Z\propto d^{-\alpha_{w}}$, where $\alpha_{w}$ depends on the wave physics. Table 1 summarizes the principal cases.
@@ 2.5 Topological selection of binary branching
where $r_{g}^{*}=r^{*}(Q_{0}/N^{g})$ are the generation-by-generation optimal radii. This decomposition holds to $\mathcal{O}(10^{-2})$: since $\alpha_{\mathrm{local}}$ varies only across $[2.898,2.907]$, the first-order gradient term $\Phi^{\prime}(r_{g}^{*})(\alpha_{t}-\alpha_{\mathrm{local}})$ is a negligible correction, leaving the Hessian $\Phi^{\prime\prime}(r_{g}^{*})$ as the dominant contribution. Consequently, $\kappa_{\mathrm{eff}}(N)=k_{t}^{\mathrm{net}}(N)/k_{w}^{\mathrm{net}}(N)$ is a strictly decreasing function of NN, uniquely maximized by binary bifurcation ($N=2$).
@@
Numerical evaluation for porcine coronary parameters ($M=2^{11}$ terminals, full three-term cost of Eq. (4)) gives the topological hierarchy shown in Table 4. The topological selection operates not by exclusion—all values $N\leq 6$ fall within the morphometric range $\alpha_{\mathrm{exp}}=2.70\pm 0.20$—but by robustness maximization. Binary bifurcation ($N=2$) uniquely places the minimax optimum at the upper end of the distribution, maximizing the separation from the wave attractor $\alpha_{w}=2.115$ and thereby achieving the greatest possible margin of transport efficiency. The stiffness ratio $\kappa_{\mathrm{eff}}$ decreases strictly with NN (from $4.98$ at $N=2$ to $1.35$ at $N=6$), monotonically reducing the wave-transport equilibrium separation. The universal biological preference for binary branching is therefore a rigorous consequence of dynamic robustness optimization: $N=2$ is the unique topology that maximizes $\kappa_{\mathrm{eff}}$, and hence the unique topology that keeps the wave-transport minimax equilibrium as far as possible from the wave-dominated limit. However, the shallow sensitivity of $\alpha^{*}$ to the branching number (varying by only $\sim\!1\%$ between $N=2$ and $N=4$ in Table 4) implies a quasi-degenerate thermodynamic landscape. Rather than a model weakness, this shallow gradient explains the biological reality that occasional trifurcations ($N=3$) are not catastrophic failures of the variational principle, but sub-optimal yet accessible morphological solutions lying only slightly above the global $\kappa_{\text{eff}}$-robustness maximum.
@@ 3.2 Evolutionary Selection via Worst-Case Robustness
The optimal duty cycle $\eta^{*}$ determined by Eq. (18) represents an architectural constraint of the network architecture. Since the transport gradient is 55 times steeper than the wave gradient at the saddle point ($|g^{\prime}|/f^{\prime}\approx 5$), the value $\eta^{*}=5/(5+1)\approx 0.833$ is strictly required for the zero-sum game to reach equilibrium. This identifies the duty cycle not as a physiological variable, but as a structural invariant of binary cardiovascular networks.
@@
The wave cost $f(\alpha)=1-(1-|\Gamma(\alpha)|^{2})^{G}$ and its derivative $f^{\prime}(\alpha^{*})$ depend only on $\alpha_{w}$, $G$, and NN (Eq. (2)). The transport cost $g(\alpha)$ defined by Eq. (6) is a ratio of sums of $\Phi$; its derivative $g^{\prime}(\alpha^{*})$ is likewise a ratio in which the metabolic prefactors ($b$, $m_{w}$, $\mu$) appear identically in every numerator and denominator term and therefore cancel exactly. The residual dependence on $Q_{0}$ and $\ell_{0}$ also cancels by the intensivity established in Proposition 7. Hence $\eta^{*}=R/(1+R)$ with $R=|g^{\prime}|/f^{\prime}$ depends only on $(G,p,\alpha_{w})$, confirming the claim. The sensitivity table (Table 7) provides numerical confirmation: all metabolic parameters yield $|S_{x}|<0.01$ for $\alpha^{*}$ and therefore also for $\eta^{*}$. ∎
@@ 4.1 Numerical inputs
As shown in Table 7, the minimax prediction $\alpha^{*}=2.72$ exhibits remarkable structural robustness. The exponent is essentially invariant ($|S_{x}|<0.01$) to metabolic ’noise’ such as absolute blood viscosity, proximal flow rate, or specific wall metabolic rates. It is determined strictly by the network topology ($G$) and the histological scaling ($p$), confirming it as a fundamental architectural property rather than a physiological tuning. Notably, the prediction is structurally independent of the absolute magnitude of the wall pre-factor $c_{0}$, as the scaling penalty depends only on the relative power pp.
@@ 4.2 Central result
The Lagrange multiplier of the equal-cost constraint evaluates to $\eta^{*}=0.833$, implying that at the saddle point the transport cost gradient is steeper than the wave cost gradient (ratio $5:1$). This apparent inversion—where the static metabolic gradient outweighs the dynamic wave sensitivity—arises from the super-linearity of the radius deviation. When the global branching exponent $\alpha$ deviates from the morphometric optimum, the metabolic cost of the millions of distal vessels explodes faster than the junction-level wave reflection. This ratio implies that the coronary vasculature sits at an architecture where the transport cost is locally 55 times more sensitive to $\alpha$ perturbations than wave costs. The distinguishing structural property of binary branching is precisely that $N=2$ maximizes the network stiffness ratio $\kappa_{\mathrm{eff}}$, providing the greatest separation between these competing gradients. As a plausibility check, the pulsatility index of coronary flow ($\mathrm{PI}=(Q_{\max}-Q_{\min})/Q_{\mathrm{mean}}\approx 2$44 [13]) independently confirms a strongly pulsatile regime in which wave-reflection costs dominate over cycle-averaged transport costs, consistent with $\eta^{*}>0.5$.
@@ 6.3 Testable predictions
The $40\%$ variation of $k_{t}^{\mathrm{net}}$ across $N\in[2,6]$ (Table 4) arises from a competition between two opposing effects: (a) the tree becomes shallower as NN grows, reducing the total number of summation terms and suppressing the depth-squared weighting; (b) the distal vessels in a high-NN tree carry lower flow, placing them in the regime where $\alpha^{*}(Q)\to(5+p)/2$ (Corollary 7 of the companion paper [10]), where the transport curvature $\Phi^{\prime\prime}(r^{*})$ is sharper. These two effects partially cancel, producing a quasi-constant $k_{t}^{\mathrm{net}}$. The net result is that the topological competition is dominated by the wave-cost curvature $k_{w}^{\mathrm{net}}\propto\ln N$, giving $\kappa_{\mathrm{eff}}(N)\propto(\ln N)^{-2.3+1}$, a strictly decreasing function of NN.